A flask is first evacuated so that it contains no gas at all. Then, 2.2 g of CO2 is introduced into the flask. On warming to 22 °C, the gas exerts a pressure of 318 mm Hg. What is the volume of the flask? 



1
Expert's answer
2021-10-05T01:36:18-0400

P(CO2) = 318 mmHg

V(CO2) = unknown

m(CO2) = 2.2 g

R = 62.3637 L mmHg mol-1 K-1

T = 22°C = (22 + 273) = 295 K


Solution:

The molar mass of CO2 is 44 g/mol

Hence,

(2.2 g CO2) × (1 mol CO2 / 44 g CO2) = 0.05 mol CO2

n(CO2) = 0.05 mol


The ideal gas equation can be used.

The ideal gas equation can be expressed as:

PV = nRT

Rearranging the ideal gas equation gives:

V = nRT / P

Hence,

V(CO2) = (0.05 mol × 62.3637 L mmHg mol-1 K-1 × 295 K) / (318 mmHg) = 2.89 L = 2.9 L

V(CO2) = 2.9 L

V(flask) = V(CO2) = 2.9 L


Answer: The volume of the flask is 2.9 liters